OpenStudy (anonymous):

Solve equation 2/7x + 1 = 1/4 + x

OpenStudy (lgbasallote):

\[\frac 27 x + 1 = \frac 14 + x\] is that your question?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\frac{ 2 }{ 7 }x +1-1-x=\frac{ 1 }{ 4 }+x-x-1\]\[\frac{ 2 }{ 7 }x-\frac{ 7 }{ 7 }x=\frac{ 1 }{ 4 }-\frac{ 4 }{ 4 }\]\[\frac{ -5 }{ 7 }x=\frac{ -3 }{ 4 }\]\[x=\frac{ -3 }{ 4 }\div \frac{ -5 }{ 7 }\]\[x=\frac{ -3 }{ 4 }\times \frac{ 7 }{ -5 }\] then you can solve x

OpenStudy (anonymous):